# CALM-PDE: Continuous and Adaptive Convolutions for Latent Space Modeling of Time-dependent PDEs

### v1

@@ 6 Experiments and Evaluation
6.1 Datasets
For regularly sampled spatial domains, we conduct experiments on the 1D Burgers’ equation dataset from PDEBench (Takamoto et al., 2022), the 2D Navier-Stokes equation datasets introduced by Li et al. (2020a), and the 3D compressible Navier-Stokes dataset of Takamoto et al. (2022). The 2D Euler equation with an airfoil geometry and the incompressible Navier-Stokes equation with cylinder geometries datasets from Pfaff et al. (2021) are used to evaluate the models on irregularly sampled domains. We refer to Appendix H for additional details on the datasets.
6.2 Baseline Models
We compare the CALM-PDE model against FNO (Li et al., 2020a) and Geo-FNO (Li et al., 2023c) for irregularly sampled meshes, OFormer (Li et al., 2023a), PIT (Chen & Wu, 2024), and LNO (Wang & Wang, 2024). We train the autoregressive models (FNO, Geo-FNO, and PIT) with a curriculum strategy that slowly increases the trajectory length which improves the error compared to training with a full autoregressive rollout (see Appendix I). OFormer uses a similar strategy, and LNO uses one-step training as proposed by the authors.
6.3 Results
We report the mean values and percentage deviations of the relative L2 error of multiple runs with different initializations. We opt for the relative L2 error as an evaluation metric because it weights channels with small and large magnitudes equally and does not ignore time-dependent decay effects like in the 1D Burgers’ equation. We refer to Section M.1 for the full results with standard deviations and Section M.2 for qualitative results.
RQ1.
We train and evaluate the models on regularly sampled spatial domains. Regularly sampled spatial domains refer to meshes with equidistant nodes or points. We select datasets that span the entire spatial dimension from 1D to 3D. Table 1 shows the relative L2 errors of the baselines and CALM-PDE model. On 2 out of 4 benchmark problems, CALM-PDE outperforms all baselines. On one problem, CALM-PDE achieves the second-lowest errors and on the remaining benchmark problem, CALM-PDE is outperformed by FNO and LNO, and outperforms PIT and OFormer.
@@ 6.3 Results
CALM-PDE is also designed to support irregularly sampled spatial domains. Irregularly sampled domains are characterized by a mesh with varying distances between the nodes. Geometries mainly cause this irregularity because the mesh is usually denser in regions of interest, such as the tip of an airfoil, where a higher accuracy is required. We train and evaluate the models on two datasets, namely the airfoil and cylinder datasets. The first dataset simulates the airflow around a static airfoil geometry. The geometry can be considered static because it does not change between different training and test samples. In the second dataset, the water flow in a channel with a cylinder as an obstacle is simulated. In contrast to the airfoil geometry, the cylinder geometry changes, which means that each training and test sample has a different cylinder geometry with a different diameter and position. Section 6 shows the relative L2 errors for both benchmark problems. On the airfoil dataset, Geo-FNO outperforms all models, while our model outperforms the Transformer-based methods, and on the cylinder dataset, CALM-PDE outperforms all baselines. Thus, CALM-PDE not only supports irregularly sampled domains but also generalizes across different geometries.
@@
We measure the inference times on an NVIDIA A100 GPU to evaluate the efficiency of CALM-PDE. Section 6.3 shows the inference times measured on the 2D Navier-Stokes dataset with a resolution of $64\times 64$ and 200 trajectories. We increase the trajectory length to measure the scaling behavior. CALM-PDE is significantly faster than OFormer and LNO and achieves a competitive inference time to that of FNO and PIT, which are fast baselines due to the use of the Fast Fourier Transform (FFT) and position attention, respectively. Section 6.3 shows the time per epoch and memory consumption for the forward and backward pass during training on 2D Navier-Stokes with a batch size of 32. CALM-PDE outperforms OFormer and LNO in terms of time and memory consumption.
@@
RQ4.
Further, we investigate the learned positions of the query points and the represented information (local or global information). The results show that the model places more query points in regions that can intuitively be considered important and that the tokens primarily learn local information. Section 6 illustrates the learned query positions for the decoder on the cylinder dataset, showing that the model samples more query points in areas where the cylinders are located. We refer to Appendix K for more details. Thus, learnable query points allow the model a higher information density in important regions and allow learning and discovering unknown important regions.
RQ5.
Finally, we analyze the impact of the model’s components in the ablation study presented in Appendix L. The results indicate that learnable query points reduce the error more effectively than both fixed and randomly sampled query points, as well as fixed query points sampled from the underlying mesh. Additionally, the kernel modulation improves the error for regularly and irregularly sampled spatial domains. Furthermore, the distance weighting, which is part of the locality inductive bias, also helps the model to further reduce the error and stabilize the training.
7 Limitations
@@ 8 Conclusion and Future Work
With CALM-PDE, we propose an efficient framework for reduced-order modeling of arbitrarily discretized PDEs. The experiments demonstrate that CALM-PDE achieves low errors on regularly and irregularly discretized PDEs, as well as that the model can generalize to different geometries. Beyond PDEs, CALM layers have broader applicability, including potential use in fields like chemistry. For future work, we aim to investigate “dynamic query points” that depend on the IC and geometry or move during the rollout to track important regions. Additionally, future work could explore improving the model for a two-stage training procedure.
@@
Funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2075 – 390740016. We acknowledge the support of the Stuttgart Center for Simulation Science (SimTech). The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Daniel Musekamp and Mathias Niepert. Additionally, we acknowledge the support of the German Federal Ministry of Education and Research (BMBF) as part of InnoPhase (funding code: 02NUK078). Lastly, we acknowledge the support of the European Laboratory for Learning and Intelligent Systems (ELLIS) Unit Stuttgart.
@@ Appendix B Details on the Problem Definition
We use discretized data generated by numerical solvers to train and test the surrogate models. The temporal domain $[0,T]$ is discretized into $N_{t}$ timesteps yielding a sequence $(u^{0},u^{t_{1}},\ldots,u^{T})$ with $N_{t}$ elements which describes the evolution of the PDE. $\Delta t:=t_{i+1}-t_{i}$ denotes the temporal step size or resolution. Similarly, the spatial domain $\Omega$ is also discretized into a finite set of N points $\{\omega_{n}\}_{n=1}^{N}$ by discretizing each spatial dimension which yields a discretized representation $\{u^{t}(\omega_{n})\}_{n=1}^{N}$ of the function. Figure 3 visualizes the discretization process. A dataset $\mathcal{D}=\{(\boldsymbol{X_{1}},\boldsymbol{Y_{1}}),\ldots,(\boldsymbol{X_{N_{s}}},\boldsymbol{Y_{N_{s}}})\}$ for each PDE consists of $N_{s}$ samples. $\boldsymbol{X_{j}}=\{u^{0}(\omega_{n})\}_{n=1}^{N}$ denotes the IC and $\boldsymbol{Y_{j}}=(\{u^{t_{1}}(\omega_{n})\}_{n=1}^{N},\ldots,\{u^{T}(\omega_{n})\}_{n=1}^{N})$ denotes the target sequence of timesteps.
@@ Appendix D Deep Parametric Continuous Convolutional Neural Networks
Figure 4 visualizes continuous convolution with and without a receptive field for 4 input points $\alpha_{n}$ and one query point $a_{1}$. As proposed by Wang et al. (2018), the kernel function $k$ is parametrized by an MLP. However, continuous convolution can be optimized further for the encoding and decoding of PDE solutions. For instance, CALM layers learn the position of the query points $a_{j}$ to effectively sample the spatial domain, use a locality inductive bias including an explicit weighting with the Euclidean distance to emphasize the input points closer to the query point and an epsilon neighborhood to improve efficiency, and each query point modulates the MLP to support different kernels for query points that have similar translation vectors.
@@ E.1 CALM Layer
Figure 5 shows a CALM layer that performs continuous convolution with three distinct features: (i) locality inductive bias, consisting of a limited receptive field and distance weighting, to emphasize closer points, (ii) learnable query points to let the model learn where to compute convolution, and (iii) kernel modulation to allow different query points to have different weights for similar translation vectors. Depending on the purpose (e.g., encoder or final layer of the decoder), the query points can be learned or provided as an external input query. The kernel function $k$ is given as
@@
The receptive field of a query point $a_{j}$ is given as $\mathtt{RF}(a_{j})=\{\alpha_{n}\in\mathcal{A}\mid\left\lVert a_{j}-\alpha_{n}\right\rVert\leq\epsilon(a_{j})\}$ where $\epsilon(a_{j})$ is the $p^{th}$ percentile of the Euclidean distances from query point $a_{j}$ to all input points $\alpha_{n}\in\mathcal{A}$ similar to Chen & Wu (2024). The hyperparameter $p$ controls the size of the receptive field. For instance, $p=0.1$ means that the epsilon neighborhoods are constructed so that each query point $a_{j}$ aggregates at least $10$% of all input points. Consequently, the epsilon neighborhood is smaller in densely sampled regions than in sparsely sampled regions which avoids convolution or aggregation with too many or few input points.
@@ E.2 Neural Architecture
In this section, we provide additional details on the neural architecture of CALM-PDE. Figure 6 shows the architecture of the model with the encoder, processor, and decoder.
@@ Appendix F Continuation of Related Work
Discretization-agnostic architectures for PDE-solving decouple the input and output spatial domains and allow solving arbitrarily discretized PDEs. Transformer-based models such as OFormer, UPT, AROMA, and LNO are discretization-agnostic. Variants such as Geo-FNO (Li et al., 2023c) of the Fourier Neural Operator (Li et al., 2020a) are also discretization-agnostic and support arbitrarily discretized inputs. Neural fields are also suitable for discretization-agnostic models and have been used in DINo (Yin et al., 2023), CORAL (Serrano et al., 2023), and Equivariant Neural Fields (Knigge et al., 2024) for PDE-solving. CALM-PDE is based on continuous convolution and also supports the encoding and decoding of arbitrarily discretized and irregularly sampled solutions. Furthermore, CALM-PDE decouples the input from the output domain.
@@ Appendix G Comparison to Related Models
Table 2 provides an overview of similarities and differences between related work and the CALM-PDE framework. We compare the following properties:
@@
Attention or Kernel Weights solely based on Positions
@@ Appendix H Additional Details on the Datasets
We benchmark the baselines and CALM-PDE model on the following datasets with regularly and irregularly sampled spatial domains. We use the term regularly sampled spatial domain to refer to spatial points sampled from a uniform grid, in contrast to the term irregularly sampled domain, which refers to a grid or mesh with non-equidistant points. Meshes are mainly irregular due to geometries or obstacles in the fluid flow. Table 3 shows an overview of the used datasets.
@@ H.1 Regularly Sampled Spatial Domains
where $v=\nabla\times u$ denotes the vorticity, $u$ the velocity, and $\nu\in\mathbb{R}_{+}$ the viscosity coefficient. $v_{0}$ is the initial vorticity and $f$ is the forcing term. The task is to learn a neural surrogate to approximate the function $v$. We use the dataset proposed by Li et al. (2020a) with periodic boundary conditions and consider the viscosities $\nu=1e^{-4}$ and $\nu=1e^{-5}$. We drop the first 10 timesteps due to less complex dynamics and consider a temporal horizon of $t\in(10,30]$ (20 timesteps) for $\nu=1e^{-4}$ and use a temporal horizon of $t\in(0,20]$ (20 timesteps) for $\nu=1e^{-5}$. The spatial domain is discretized into a grid of $64\times 64$. We use the last 200 samples for training and the remaining ones for training in both cases.
@@
where $\rho$ is the density, $\bf{v}$ the velocity, and $p$ the pressure of the fluid. $\epsilon$ denotes the internal energy and $\sigma^{\prime}$ the viscous stress tensor. The parameters $\eta$ and $\zeta$ are the shear and bulk viscosity. Equation 20a represents the conservation of mass, Equation 20b is the equation of the conservation of momentum, and Equation 20c is the energy conservation. The goal is to learn a neural surrogate approximating density, velocity, and pressure. We use the dataset from PDEBench (Takamoto et al., 2023) with $\eta=\zeta=1e^{-8}$, a spatial resolution of $64\times 64\times 64$, and the full temporal resolution of 21 timesteps. The equation in the dataset is under periodic boundary conditions. We use the first 10 samples for testing and the last 90 samples for training.
@@ H.2 Irregularly Sampled Spatial Domains
$\rho:=\rho(t,\boldsymbol{\omega}),\mathbf{v}:=\mathbf{v}(t,\boldsymbol{\omega})$ $,p:=p(t,\boldsymbol{\omega}),\qquad\boldsymbol{\omega}\in\Omega,t\in[0,T]$
@@
$\mathbf{v}:=\mathbf{v}(t,\boldsymbol{\omega})$ $,p:=p(t,\boldsymbol{\omega}),\qquad\boldsymbol{\omega}\in\Omega,t\in[0,T]$

### v2

@@ 6.3 Results
CALM-PDE is also designed to support irregularly sampled spatial domains. Irregularly sampled domains are characterized by a mesh with varying distances between the nodes. Geometries mainly cause this irregularity because the mesh is usually denser in regions of interest, such as the tip of an airfoil, where a higher accuracy is required. We train and evaluate the models on two datasets, namely the airfoil and cylinder datasets. The first dataset simulates the airflow around a static airfoil geometry. The geometry can be considered static because it does not change between different training and test samples. In the second dataset, the water flow in a channel with a cylinder as an obstacle is simulated. In contrast to the airfoil geometry, the cylinder geometry changes, which means that each training and test sample has a different cylinder geometry with a different diameter and position. Section˜6.3 shows the relative L2 errors for both benchmark problems. On the airfoil dataset, AROMA achieves the lowest error overall, while our model outperforms Transformer-based baselines such as OFormer and LNO, and on the cylinder dataset, CALM-PDE delivers the second-best performance, closely trailing AROMA. Thus, CALM-PDE not only supports irregularly sampled domains but also generalizes across different geometries.
@@
We measure the inference times on an NVIDIA A100 GPU to evaluate the efficiency of CALM-PDE. Section˜7 shows the inference times measured on the 2D Navier-Stokes dataset with a resolution of $64\times 64$ and 200 trajectories. We increase the trajectory length to measure the scaling behavior. CALM-PDE is significantly faster than OFormer and LNO and achieves a competitive inference time to that of FNO and PIT, which are fast baselines due to the use of the Fast Fourier Transform (FFT) and position attention, respectively. Section˜7 shows the time per epoch and memory consumption for the forward and backward pass during training on 2D Navier-Stokes with a batch size of 32. CALM-PDE outperforms OFormer and LNO in terms of time and memory consumption.
RQ4.
Further, we investigate the learned positions of the query points and the represented information (local or global information). The results show that the model places more query points in regions that can intuitively be considered important and that the tokens primarily learn local information. Section˜6.3 illustrates the learned query positions for the decoder on the cylinder dataset, showing that the model samples more query points in areas where the cylinders are located. We refer to Appendix˜L for more details. Thus, learnable query points allow the model a higher information density in important regions and allow learning and discovering unknown important regions.
RQ5.
Finally, we analyze the impact of the model’s components in the ablation study presented in Appendix˜M. The results indicate that learnable query points reduce the error more effectively than both fixed and randomly sampled query points, as well as fixed query points sampled from the underlying mesh. Additionally, the kernel modulation improves the error for regularly and irregularly sampled spatial domains. Furthermore, the distance weighting, which is part of the locality inductive bias, also helps the model to further reduce the error and stabilize the training.
7 Limitations
@@ 8 Conclusion and Future Work
With CALM-PDE, we propose an efficient framework for reduced-order modeling of arbitrarily discretized PDEs. The experiments demonstrate that CALM-PDE achieves low errors on regularly and irregularly discretized PDEs, as well as that the model can generalize to different geometries. While it does not consistently surpass all baselines, it delivers competitive errors across various problems. This makes it a viable alternative to established approaches such as FNOs and transformer-based architectures and demonstrates that convolution can also be applied to irregular meshes. Beyond PDEs, CALM layers have broader applicability, including potential use in fields like chemistry. For future work, we aim to improve CALM-PDE for problems with high frequencies, which are important in real-world applications. Additionally, future work could investigate “dynamic query points” that depend on the IC and geometry or move during the rollout to track important regions. Another interesting direction for future research is improving the model for a two-stage training procedure.
@@
Funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2075 – 390740016. We acknowledge the support of the Stuttgart Center for Simulation Science (SimTech). The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Daniel Musekamp and Mathias Niepert. Additionally, we acknowledge the support of the German Federal Ministry of Research, Technology and Space (BMFTR) as part of InnoPhase (funding code: 02NUK078). Lastly, we acknowledge the support of the European Laboratory for Learning and Intelligent Systems (ELLIS) Unit Stuttgart.
@@ Appendix B Details on the Problem Definition
We use discretized data generated by numerical solvers to train and test the surrogate models. The temporal domain $[0,T]$ is discretized into $N_{t}$ timesteps yielding a sequence $(u^{0},u^{t_{1}},\ldots,u^{T})$ with $N_{t}$ elements which describes the evolution of the PDE. $\Delta t:=t_{i+1}-t_{i}$ denotes the temporal step size or resolution. Similarly, the spatial domain $\Omega$ is also discretized into a finite set of N points $\{\omega_{n}\}_{n=1}^{N}$ by discretizing each spatial dimension which yields a discretized representation $\{u^{t}(\omega_{n})\}_{n=1}^{N}$ of the function. Figure˜3 visualizes the discretization process. A dataset $\mathcal{D}=\{(\boldsymbol{X_{1}},\boldsymbol{Y_{1}}),\ldots,(\boldsymbol{X_{N_{s}}},\boldsymbol{Y_{N_{s}}})\}$ for each PDE consists of $N_{s}$ samples. $\boldsymbol{X_{j}}=\{u^{0}(\omega_{n})\}_{n=1}^{N}$ denotes the IC and $\boldsymbol{Y_{j}}=(\{u^{t_{1}}(\omega_{n})\}_{n=1}^{N},\ldots,\{u^{T}(\omega_{n})\}_{n=1}^{N})$ denotes the target sequence of timesteps.
@@ Appendix D Deep Parametric Continuous Convolutional Neural Networks
Figure˜4 visualizes continuous convolution with and without a receptive field for 4 input points $\alpha_{n}$ and one query point $a_{1}$. As proposed by Wang et al. (2018), the kernel function $k$ is parametrized by an MLP. However, continuous convolution can be optimized further for the encoding and decoding of PDE solutions. For instance, CALM layers learn the position of the query points $a_{j}$ to effectively sample the spatial domain, use a locality inductive bias including an explicit weighting with the Euclidean distance to emphasize the input points closer to the query point and an epsilon neighborhood to improve efficiency, and each query point modulates the MLP to support different kernels for query points that have similar translation vectors.
@@ E.1 CALM Layer
Figure˜5 shows a CALM layer that performs continuous convolution with three distinct features: (i) locality inductive bias, consisting of a limited receptive field and distance weighting, to emphasize closer points, (ii) learnable query points to let the model learn where to compute convolution, and (iii) kernel modulation to allow different query points to have different weights for similar translation vectors. Depending on the purpose (e.g., encoder or final layer of the decoder), the query points can be learned or provided as an external input query. The kernel function $k$ is given as
@@
The receptive field of a query point $a_{j}$ is given as $\mathtt{RF}(a_{j})=\{\alpha_{n}\in\mathcal{A}\mid\left\lVert a_{j}-\alpha_{n}\right\rVert\leq\epsilon(a_{j})\}$ where $\epsilon(a_{j})$ is the $p^{th}$ percentile of the Euclidean distances from query point $a_{j}$ to all input points $\alpha_{n}\in\mathcal{A}$ similar to Chen & Wu (2024). The hyperparameter $p$ controls the size of the receptive field. For instance, $p=0.1$ means that the epsilon neighborhoods are constructed so that each query point $a_{j}$ aggregates at least 1010% of all input points. Consequently, the epsilon neighborhood is smaller in densely sampled regions than in sparsely sampled regions which avoids convolution or aggregation with too many or few input points.
@@
Computational Complexity.
The complexity of a CALM layer is given as follows. Let $N$ be an arbitrary number of input points. The output of the layer are features for $N^{\prime}$ output or query points, where $N^{\prime}$ is constant for all layers except for the final decoder layer. The percentile, which determines the size of the receptive field, and the feature or channel dimension of the CALM layer are denoted as $p$ and $d$, respectively. This results in $\mathcal{O}(p\cdot N\cdot N^{\prime}\cdot d)$ computations in total. In contrast, models such as AROMA (Serrano et al., 2024) or LNO (Wang & Wang, 2024) do not consider a receptive field, which requires $\mathcal{O}(N\cdot N^{\prime}\cdot d)$ computations for $N^{\prime}$ query tokens. Since $p$ is usually small (e.g., $p=0.01$), the complexity is reduced by two orders of magnitude compared to models without a receptive field.
@@ E.2 Neural Architecture
In this section, we provide additional details on the neural architecture of CALM-PDE. Figure˜6 shows the architecture of the model with the encoder, processor, and decoder.
@@ Appendix F Continuation of Related Work
Discretization-agnostic architectures for PDE-solving decouple the input and output spatial domains and allow solving arbitrarily discretized PDEs. Transformer-based models such as OFormer, UPT, AROMA, and LNO are discretization-agnostic. Variants such as Geo-FNO (Li et al., 2023c) of the Fourier Neural Operator (Li et al., 2021b) are also discretization-agnostic and support arbitrarily discretized inputs. Neural fields are also suitable for discretization-agnostic models and have been used in DINo (Yin et al., 2023), CORAL (Serrano et al., 2023), and Equivariant Neural Fields (Knigge et al., 2024) for PDE-solving. CALM-PDE is based on continuous convolution and also supports the encoding and decoding of arbitrarily discretized and irregularly sampled solutions. Furthermore, CALM-PDE decouples the input from the output domain.
@@ Appendix G Comparison to Related Models
Table˜2 provides an overview of similarities and differences between related work and the CALM-PDE framework. We compare the following properties:
@@
Attention or Kernel Weights
solely based on Positions
@@ Appendix H Additional Details on the Datasets
We benchmark the baselines and CALM-PDE model on the following datasets with regularly and irregularly sampled spatial domains. We use the term regularly sampled spatial domain to refer to spatial points sampled from a uniform grid, in contrast to the term irregularly sampled domain, which refers to a grid or mesh with non-equidistant points. Meshes are mainly irregular due to geometries or obstacles in the fluid flow. Table˜3 shows an overview of the used datasets.
@@ H.1 Regularly Sampled Spatial Domains
where $v=\nabla\times u$ denotes the vorticity, $u$ the velocity, and $\nu\in\mathbb{R}_{+}$ the viscosity coefficient. $v_{0}$ is the initial vorticity and $f$ is the forcing term. The task is to learn a neural surrogate to approximate the function $v$. We use the dataset proposed by Li et al. (2021b) with periodic boundary conditions and consider the viscosities $\nu=1e^{-4}$ and $\nu=1e^{-5}$. We drop the first 10 timesteps due to less complex dynamics and consider a temporal horizon of $t\in(10,30]$ (20 timesteps) for $\nu=1e^{-4}$ and use a temporal horizon of $t\in(0,20]$ (20 timesteps) for $\nu=1e^{-5}$. The spatial domain is discretized into a grid of $64\times 64$. We use the last 200 samples for testing and the remaining ones for training in both cases.
@@
where $\rho$ is the density, $\bf{v}$ the velocity, and $p$ the pressure of the fluid. $\epsilon$ denotes the internal energy and $\sigma^{\prime}$ the viscous stress tensor. The parameters $\eta$ and $\zeta$ are the shear and bulk viscosity. Equation˜20a represents the conservation of mass, Equation˜20b is the equation of the conservation of momentum, and Equation˜20c is the energy conservation. The goal is to learn a neural surrogate approximating density, velocity, and pressure. We use the dataset from PDEBench (Takamoto et al., 2023) with $\eta=\zeta=1e^{-8}$, a spatial resolution of $64\times 64\times 64$, and the full temporal resolution of 21 timesteps. The equation in the dataset is under periodic boundary conditions. We use the first 10 samples for testing and the last 90 samples for training.
@@ H.2 Irregularly Sampled Spatial Domains
$\rho=\rho(t,\boldsymbol{\omega}),\mathbf{v}=\mathbf{v}(t,\boldsymbol{\omega})$ $,p=p(t,\boldsymbol{\omega}),\qquad\boldsymbol{\omega}\in\Omega,t\in[0,T]$
@@
$\mathbf{v}=\mathbf{v}(t,\boldsymbol{\omega})$ $,p=p(t,\boldsymbol{\omega}),\qquad\boldsymbol{\omega}\in\Omega,t\in[0,T]$